OpenStudy (anonymous):

A bag contains 50 marbles, 10 of which are blue, 8 are red, 20 are green, and 12 are purple. Marcie takes a marble out of the bag, records the color, and returns it to the bag. How many red marbles should she expect after 500 trials?

OpenStudy (anonymous):

@tkhunny

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

are there any answer choices?

OpenStudy (anonymous):

80 100 120 200

OpenStudy (anonymous):

not completely sure but i think its 120

OpenStudy (phi):

Do you know how to find the probability of picking a red marble?

OpenStudy (anonymous):

no

OpenStudy (phi):

Probability is # of "good" choices / total number in this case # of red / total number of marbles what do you get ?

OpenStudy (anonymous):

50 /8 6.25 but 500/8 62.5

OpenStudy (amistre64):

the ratio of 50 to 8 remains constant: 50n to 8n, such that 50n=500, what is n?

OpenStudy (tkhunny):

10 of which are blue, 8 are red, 20 are green, and 12 are purple Checking: 10 + 8 + 20 + 12 = 50 -- Good Theoretical Ratios: Blue 10/50 = 20%, Red 8/50 = 16%, Green 20/50 = 40%, Purple 12/50 = 24% Checking: 20 + 16 + 40 + 24 = 100 -- Good! See how that works? Mind you, this is WITH REPLACEMENT. The probabilities and ratios are constant. If you start throwing them away, rather than putting them back, it's a different world.